For the production function y = KαL β , where α, β ∈ (0, 1)

(a) Derive the marginal production of K and L. Show that it is decreasing.


1
Expert's answer
2021-11-04T16:21:35-0400

For the production function y = KαL β , where α, β ∈ (0, 1)


Marginal product of K( MP"_{K}") is;


"\frac{\delta y}{\delta K}=\alpha K^{\alpha-1}L^{\beta}"


Marginal product of L (MP"_{L}" ) is:


"\frac{\delta y}{\delta L}=\beta K^{\alpha}L^{\beta-1}"


To show that the marginal product of capital diminishes as the capital input increases with the labour input held constant, I calculate its derivative with respect to K:



"\frac{\delta MP_{K}}{\delta K}=" "\alpha^{2}-\alpha K^{\alpha-2}L^{\beta}"


To show that the marginal product of labour decreases as the labour input increases with the capital input held constant, we calculate its derivative with respect to L:


"\frac{\delta MP_{L}}{\delta L}=" "\beta^{2}-\beta K^{\alpha}L^{\beta-2}"